How would you write an equation for this word problem?
-A student has a part time job at an electronics store. She earns $85 for working 12h each week, plus 5% of her sales over $1000 each week.
I need an equation and then to find the function g(5000). What i've tried is g(x)=85+0.05(x-1000) but Im not sure if it would work. I can't remember how to incorporate the percentage in the equation because wouldn't there need to be 85x since it's $85 for every 12 hours if x=hours and y=salary

Answer by stanbon(75887)

-A student has a part time job at an electronics store. She earns $85 for working 12h each week, plus 5% of her sales over $1000 each week.
I need an equation and then to find the function g(5000).
----------------------------
g(x hrs) = 85 + 0.05(x-1000)
---------------------------------------
g(5000) = 85 + 0.05*(5000-1000)-50
------
g(5000) = 85 + 0.05*4000 - 50
------
g(5000) = 85 + 200 - 50 = $235.00
=========================
Cheers,
Stan H.
========================